simplify:
(4x^3+x^2-3) - (-9x^3+5x^2-6)
: simplify:
(4x^3+x^2-3) - (-9x^3+5x^2-6)

Answer by Fombitz(1755)

The key is to group like terms and perform operations on them.
Group the x^2 terms, the x terms and the constants.
Use the distributive properties and watch the signs.
(4x^3+x^2-3) - (-9x^3+5x^2-6)=(4x^3-(-9x^3))+(x^2-5x^2)+(-3-(-6))
(4x^3+x^2-3) - (-9x^3+5x^2-6)=(4x^3+9x^3)+(x^2-5x^2)+(-3+6)
(4x^3+x^2-3) - (-9x^3+5x^2-6)=(13x^3)+(-4x^2)+(3)
(4x^3+x^2-3) - (-9x^3+5x^2-6)=13x^3-4x^2+3